OpenStudy (anonymous):

Find a third-degree polynomial function such that f(0) = 18 and whose zeros are -1, 2, and 3. Using complete sentences, explain how you found it.

OpenStudy (anonymous):

Since the zeros of the third degree polynomial are -1,2 and 3 f(x) = k (x+1)(x-2)(x-3) Now, f(0)=18 k(0+1) (0-2) (0-3) =18 k=3 So, f(x) =3(x+1)(x-2)(x-3)

OpenStudy (anonymous):

got it?